# DU-Shapley: A Shapley Value Proxy for Efficient Dataset Valuation

### v2

@@ Appendix B Complementary Numerical results
B.2 Approximating the Shapley value in Synthetic Data
@@ B.3 Owen’s Shapley value approximation
In Section B.2, we considered the Shapley value approximation referred to as Owen-Shapley as a state-of-the-art competitor to DU-Shapley. We provide in the following additional details regarding Owen-Shapley. For the other competitors, we directly refer the interested reader to Mitchell et al. [30]. Owen [32] studied the multilinear extension of a cooperative game and an alternative way to express the Shapley value. Formally, a cooperative game $G=(\mathcal{I},u)$ consists on a set of $I$ players =$\mathcal{I}=\{1,2,...,I\}$ and a value function $u:2^{\mathcal{I}}\to\mathbb{R}$ such that, for any $S\subseteq\mathcal{I}$, $u(S)$ corresponds to the value generated by the coalition $S$. The multilinear extension of $G$, denoted $\bar{G}=(\mathcal{I},\bar{u})$, is obtained when considering the value function $\bar{u}:[0,1]^{\mathcal{I}}\to\mathbb{R}$ given by,
@@
B.4 OpenDataVal implementations
$\mathrm{LOO}_{i}:=u(\mathcal{I})-u(\mathcal{I}\setminus\{i\}).$ (27)
As players’ marginal contributions to large datasets tend to be small, we have preferred to sample one dataset $\mathrm{D}$ from $\mathrm{D}_{i}$ and to output
$\mathrm{LOO}_{i}:=u(\mathrm{D}\cup\mathrm{D}_{i})-u(\mathrm{D}).$ (28)
@@ Appendix C Missing proofs
Since players distributions differ on their variances, $\bigl{(}\sum_{i\in\mathcal{S}}X_{i}^{\top}\varepsilon_{i}^{-2}X_{i}\bigr{)}^{-1}$ corresponds to a semi-correlated Wishart random variable whose expected value can be approximated [35] by
@@ C.2 Proof of Theorem 1
all exist, for some constants $\mu_{A},\mu_{B},\sigma_{A},\sigma_{B}>0$. Let $K\sim\mathrm{U}(\{0,\ldots,I\})$, $\mathcal{S}_{K}\sim\mathrm{U}([2^{\mathcal{I}}_{K}])$, and define $q(\mathcal{S}_{K})$ as in (14) for the third use-case. Then, almost surely, $\nicefrac{{q(\mathcal{S}_{K})}}{{q(\mathcal{I})}}\xrightarrow{I\to\infty}\mathrm{U}([0,1])$.
@@
and that $K=\lfloor IU\rfloor$ with $U\sim\mathrm{U}([0,1])$. ∎
@@ C.3.1 Technical lemmata
Let $I\in\mathbb{N}$, $N:=\{n_{1},\ldots,n_{I}\}\in\mathbb{R}_{+}^{I}$, $\mu=\frac{1}{I}\sum_{i=1}^{I}n_{i}$ be their mean value and $\sigma^{2}=\frac{1}{I}\sum_{i=1}^{I}(n_{i}-\mu)^{2}$ be their variance. For $k\in\{0,\ldots,n\}$, let $\mathcal{S}_{k}\sim\mathrm{U}(\{S_{k}\subseteq[I]:|S_{k}|=k\})$ be a uniform random variable on the subsets of $\{1,\ldots,I\}$ of size $k$, and $n_{\mathcal{S}_{k}}=\sum_{i\in\mathcal{S}_{k}}n_{i}$ be the random variable defined by the sum of the elements of $\mathcal{S}_{k}$. Let $K\sim\mathrm{U}(\{0,\ldots,I\})$ and define $\mathbf{Y}=n_{\mathcal{S}_{K}}$. Then,
$\mathbb{E}[\mathbf{Y}-\mu K\mid K=k]=0,$ (58)
$\mathbb{E}\bigl{[}\left(\mathbf{Y}-\mu K\right)^{2}\mid K=k\bigr{]}\leq k\sigma^{2}.$ (59)
@@
Consider $\mathcal{S}_{k}$, $n_{\mathcal{S}_{k}}$, $K$, and $\mathbf{Y}$ as in Lemma 2. Let $w:\mathbb{R}_{+}\to\mathbb{R}$ be a function in $\mathcal{C}^{2}$, increasing, and suppose there exists $\kappa\in\mathbb{R}_{+}$, such that,
@@ C.3.2 Proof of Theorem 2
$\varphi_{i}=\mathbb{E}[w(\mathbf{Y}_{-i}+n_{i})-w(\mathbf{Y}_{-i})],\quad\psi_{i}=\mathbb{E}[w(K\mu_{-i}+n_{i})-w(K\mu_{-i})],$ (99)
where $K\sim\mathrm{U}([I-1])$ and $\mathbf{Y}_{-i}=n_{\mathcal{S}^{(i)}_{K}}$ with $\mathcal{S}^{(i)}_{K}$ taking values on the subsets of $\mathcal{I}\setminus\{i\}$ of size $K$, (b) writing
$|\varphi_{i}-\psi_{i}|\leq$ $|\mathbb{E}[w(\mathbf{Y}+n_{i})-w(K\mu_{-i}+n_{i})]|+|\mathbb{E}[w(\mathbf{Y})-w(K\mu_{-i})]|,$ (100)

### v3

@@ Appendix A Complementary Numerical results
A.1 DU-Shapley vs SVARM
We have looked at the probability at which SVARM (Theorem 4 [22]) can ensure, after $I^{2}$ iterations (without considering the warm up as part of the budget), an error equal to DU-Shapley’s accumulated bias. We have considered the same value function than in Section 4.2 with $n_{max}\in\{2\cdot 10^{3},3\cdot 10^{3},5\cdot 10^{3},10^{4}\}$ and 100 simulations of sets of players at each time. Figure 3 shows the results. We observe how SVARM cannot ensure, with high enough probability, an approximation error equal to the one of DU-Shapley.
Figure 3: Probability that SVARM guarantees an error equal to DU-Shapley’s bias
A.2 Approximating the Shapley value in Synthetic Data
Appendix B Further details about numerical implementations
@@
B.2 OpenDataVal implementations
$\mathrm{LOO}_{i}:=u(\mathcal{I})-u(\mathcal{I}\setminus\{i\}).$ (26)
As players’ marginal contributions to large datasets tend to be small, we have preferred to sample one dataset $\mathrm{D}$ from $\mathrm{D}_{-i}$ and to output
$\mathrm{LOO}_{i}:=u(\mathrm{D}\cup\mathrm{D}_{i})-u(\mathrm{D}).$ (27)
@@ B.3 Owen’s Shapley value approximation
In Section A.2, we considered the Shapley value approximation referred to as Owen-Shapley as a state-of-the-art competitor to DU-Shapley. We provide in the following additional details regarding Owen-Shapley. For the other competitors, we directly refer the interested reader to Mitchell et al. [32]. Owen [34] studied the multilinear extension of a cooperative game and an alternative way to express the Shapley value. Formally, a cooperative game $G=(\mathcal{I},u)$ consists on a set of $I$ players =$\mathcal{I}=\{1,2,...,I\}$ and a value function $u:2^{\mathcal{I}}\to\mathbb{R}$ such that, for any $S\subseteq\mathcal{I}$, $u(S)$ corresponds to the value generated by the coalition $S$. The multilinear extension of $G$, denoted $\bar{G}=(\mathcal{I},\bar{u})$, is obtained when considering the value function $\bar{u}:[0,1]^{\mathcal{I}}\to\mathbb{R}$ given by,
@@ Appendix C Missing proofs
Since players distributions differ on their variances, $\sum_{i\in\mathcal{S}}X_{i}^{\top}\varepsilon_{i}^{-2}X_{i}$ corresponds to a semi-correlated Wishart random variable where each $\frac{1}{\varepsilon_{i}}X_{i}\sim\mathrm{N}(0,(\frac{\sigma_{i}}{\varepsilon_{i}})^{2}\mathrm{I}_{d})$. In particular, the semi-correlated Wishart random variable can be approximated by a central Wishart distribution [37, 43], whose precision depends on the homogeneity of the coefficients $\sigma_{i}/\varepsilon_{i}$ over all $i\in\mathcal{I}$, as showed in [20]. It follows,
@@ C.2 Proof of Theorem 1
all exist, for some constants $\mu_{A},\mu_{B},\sigma_{A},\sigma_{B}>0$. Let $\mathbf{K}\sim\mathrm{U}(\{0,\ldots,I\})$, $\mathcal{S}_{\mathbf{K}}\sim\mathrm{U}([2^{\mathcal{I}}_{\mathbf{K}}])$, and define $q(\mathcal{S}_{\mathbf{K}})$ as in (14) for the third use-case. Then, almost surely, $\nicefrac{{q(\mathcal{S}_{\mathbf{K}})}}{{q(\mathcal{I})}}\xrightarrow{I\to\infty}\mathrm{U}([0,1])$.
@@
and that $\mathbf{K}=\lfloor IU\rfloor$ with $U\sim\mathrm{U}([0,1])$. ∎
@@ C.3.1 Technical lemmata
Let $I\in\mathbb{N}$, $N:=\{n_{1},\ldots,n_{I}\}\in\mathbb{R}_{+}^{I}$, $\mu=\frac{1}{I}\sum_{i=1}^{I}n_{i}$ be their mean value and $\sigma^{2}=\frac{1}{I}\sum_{i=1}^{I}(n_{i}-\mu)^{2}$ be their variance. For $k\in\{0,\ldots,n\}$, let $\mathcal{S}_{k}\sim\mathrm{U}(\{S_{k}\subseteq[I]:|S_{k}|=k\})$ be a uniform random variable on the subsets of $\{1,\ldots,I\}$ of size $k$, and $n_{\mathcal{S}_{k}}=\sum_{i\in\mathcal{S}_{k}}n_{i}$ be the random variable defined by the sum of the elements of $\mathcal{S}_{k}$. Let $\mathbf{K}\sim\mathrm{U}(\{0,\ldots,I\})$ and define $\mathbf{Y}=n_{\mathcal{S}_{\mathbf{K}}}$. Then,
$\mathbb{E}[\mathbf{Y}-\mu\mathbf{K}\mid\mathbf{K}=k]=0,$ (61)
$\mathbb{E}\bigl{[}\left(\mathbf{Y}-\mu\mathbf{K}\right)^{2}\mid\mathbf{K}=k\bigr{]}\leq k\sigma^{2}.$ (62)
@@
Consider $\mathcal{S}_{\mathbf{K}}$, $n_{\mathcal{S}_{\mathbf{K}}}$, $\mathbf{K}$, and $\mathbf{Y}$ as in Lemma 2. Let $w:\mathbb{R}_{+}\to\mathbb{R}$ be a function in $\mathcal{C}^{2}$, increasing, and suppose there exists $\kappa\in\mathbb{R}_{+}$, such that,
@@ C.3.2 Proof of Theorem 2
$\varphi_{i}=\mathbb{E}[w(\mathbf{Y}_{-i}+n_{i})-w(\mathbf{Y}_{-i})],\quad\psi_{i}=\mathbb{E}[w(\mathbf{K}\mu_{-i}+n_{i})-w(\mathbf{K}\mu_{-i})],$ (102)
where $\mathbf{K}\sim\mathrm{U}([I-1])$ and $\mathbf{Y}_{-i}=n_{\mathcal{S}^{(i)}_{\mathbf{K}}}$ with $\mathcal{S}^{(i)}_{\mathbf{K}}$ taking values on the subsets of $\mathcal{I}\setminus\{i\}$ of size $\mathbf{K}$, (b) writing
$|\varphi_{i}-\psi_{i}|\leq$ $|\mathbb{E}[w(\mathbf{Y}+n_{i})-w(\mathbf{K}\mu_{-i}+n_{i})]|+|\mathbb{E}[w(\mathbf{Y})-w(\mathbf{K}\mu_{-i})]|,$ (103)